OpenStudy (anonymous):

Simplify: 5x -5/5 A. x - 15y B. x + 15y C. x + 3y D. x - 3y

OpenStudy (anonymous):

is it \[\frac{5x - 5}{5} \space or \space 5x - \frac{5}{5}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

5x 15y/5

OpenStudy (anonymous):

\[\frac{5x - 15}{5}\]??

OpenStudy (anonymous):

If thats what it is then the answer is D